### v1

# LumiSculpt: A Consistency Lighting Control Network for Video Generation

Abstract
Lighting plays a pivotal role in ensuring the naturalness of video generation, significantly influencing the aesthetic quality of the generated content. However, due to the deep coupling between lighting and the temporal features of videos, it remains challenging to disentangle and model independent and coherent lighting attributes, limiting the ability to control lighting in video generation. In this paper, inspired by the established controllable T2I models, we propose LumiSculpt, which, for the first time, enables precise and consistent lighting control in T2V generation models.LumiSculpt equips the video generation with strong interactive capabilities, allowing the input of custom lighting reference image sequences. Furthermore, the core learnable plug-and-play module of LumiSculpt facilitates remarkable control over lighting intensity, position, and trajectory in latent video diffusion models based on the advanced DiT backbone.Additionally, to effectively train LumiSculpt and address the issue of insufficient lighting data, we construct LumiHuman, a new lightweight and flexible dataset for portrait lighting of images and videos. Experimental results demonstrate that LumiSculpt achieves precise and high-quality lighting control in video generation.
* Work done during internship at Ant Group. $\dagger$ Project lead.✉ Corresponding author.
Figure 1: LumiSculpt allows user-specified lighting intensity, direction, and trajectories, with textual conditions as inputs. Only one training session is needed for LumiSculpt to generate diverse results.
1 Introduction
“If a video tells a story, then lighting is the voice that shapes its tone and mood.”
Lighting is essential for the essential of video generation, which is one of the defining factors for the overall aesthetic quality of the generated video, and is also used to convey emotions, highlight character traits, and guide the audience’s attention. Mainstream video generation methods currently employ latent diffusion models (LDMs) to achieve video generation through multi-step denoising in latent space. Research on controllable image and video generation based on LDMs supports our studies of consistent lighting control. Several methods (Ho et al., 2022b; Chen et al., 2023; Ho et al., 2022a; Wang et al., 2023a; Guo et al., 2024) have been developed to achieve relatively accurate text-controlled video generation, as well as video editing (Chai et al., 2023; Ceylan et al., 2023), customization (Wu et al., 2023), and controlling (Wang et al., 2023b; Wei et al., 2024; He et al., 2024). These works have improved the controllability, aesthetics, and usability of video generation. However, due to the deep coupling between lighting and the temporal features of videos, it is challenging to model independent and coherent lighting attributes, resulting in a lack of handy approaches to controlling lighting in videos.
The challenge of customizing lighting lies in three aspects: the lack of training data, the representation of lighting, and the mechanism of injecting lighting features without influencing other attributes. Specifically, although there are currently relighting datasets based on light stages (Debevec et al., 2000), the data format of light stages is not easily applicable in video generation scenarios. Therefore, a flexible dataset that is adaptable to text-controlled new content generation is needed. Obtaining the projection of lighting on the camera’s imaging plane requires knowledge of the lighting and the surface texture of the illuminated object (Kim et al., 2024; Ren et al., 2024; Mei et al., 2024), which cannot be satisfied in an end-to-end video generation scenario. Thus, a simpler lighting representation that is only related to lighting parameters is important. Finally, similar to most control tasks, lighting control faces the problem of the deep decoupling of lighting from other elements, such as semantics and color.
In this paper, we propose LumiHuman to solve the problem of limited training data. LumiHuman is a portrait lighting dataset that can constitut more than 220K videos of humans with known lighting parameters. This is a lightweight and flexible dataset that is not limited to specific lighting movements but is presented in freely combinable frames, laying the foundation for a more diverse range of lighting paths and combinations. We then use virtual engine rendering with known lighting parameters to obtain projections of different directional lighting on planes as a lighting representation. Additionally, we propose a consistency lighting control network, LumiSculpt, which learns an accurate plug-and-play lighting module capable of controlling the direction and movement of lighting in video generation. To solve the problem of lighting feature injection, we introduce a light control module that takes the lighting projection as input and integrates lighting control injected into the generative model layer by layer. Furthermore, to better decouple lighting and appearance, we design a decoupling loss based on a dual-branch structure, preserving diverse generative capabilities. We implemented LumiSculpt on Open-Sora (Lab & etc., 2024) to enable precise lighting control. We conducted comprehensive quantitative and qualitative evaluations. The experimental results show that LumiSculpt has achieved state-of-the-art performance in the control of text-to-video lighting, as shown in Figure 1. In summary, our main contributions are as follows:
  • We propose a portrait lighting dataset LumiHuman, which includes 65 different individuals and allows for the free combination of over 3000 lighting trajectories. LumiHuman paves the way for more lighting control in both image and video generation.
  • We introduce LumiSculpt, enabling control of lighting direction and movement trajectories in video generation. We propose a lighting representation method, a lighting injection approach, and a lighting decoupling loss for the text-to-video generation scenario, enabling diverse content generation with limited data.
  • Extensive experiments prove that LumiSculpt has achieved state-of-the-art performance. LumiSculpt can not only generate accurate lighting intensity, direction, and trajectories but also maintain coherence and content diversity.
2 Related Works
2.1 Relighting
In recent years, deep learning techniques have made significant progress in portrait relighting (Kim et al., 2024; Mei et al., 2023; 2024; Nestmeyer et al., 2020; Pandey et al., 2021; Sun et al., 2019; Wang et al., 2020; Yeh et al., 2022; Zhang et al., 2021), often relying on paired data captured by light stage systems (Debevec et al., 2000) for supervised learning. Typically, these methods require the use of high dynamic range (HDR) environmental maps as input. This process involves estimating intermediate surface properties, including normal vectors, albedo, diffuse reflectance, and specular reflection characteristics. However, the reliance on HDR environmental maps limits the practical application of these techniques in video generation scenarios. Besides, researchers also explore portrait relighting techniques that do not depend on light stage data (Hou et al., 2021; 2022; Wang et al., 2023c).
Recently, diffusion-based models have brought new research directions to the field. Ren et al. (Ren et al., 2024) propose a three-stage lighting-aware diffusion model called Relightful Harmonization, which aims to provide complex lighting coordination for foreground portraits with any background image. Zeng et al. (Zeng et al., 2024) propose a three-stage portrait relighting method using a fine diffusion model called DiLightNet, which calculates radiance cues to re-synthesize and refine the foreground object by combining the rough shape of the foreground object inferred from the preliminary image. Xing et al. (Xing et al., 2024) propose a natural image relighting method called Retinex-Diffusion, which treats the diffusion model as a black-box image renderer and strategically decomposes its energy function to be consistent with the image formation model. However, there is still a lack of methods for lighting control in text-to-video generation.
2.2 Text-to-video synthesis and controlling
Recently, several researches, such as (Ho et al., 2022b; Chen et al., 2023; Ho et al., 2022a; Wang et al., 2023a; Guo et al., 2024), have adopted diffusion models to create highly realistic video content, utilizing text as conditions in guiding the generation process. These studies focus on ensuring consistency between textual descriptions and the final video output. Addressing the issue of difficulty in precisely describing specific visual attributes through text conditions, some studies have attempted to achieve finer video control by fine-tuning models or introducing additional control parameters. Tune-A-Video (Wu et al., 2023) propose a fine-tuning framework that allows users to customize specific videos. VideoComposer (Wang et al., 2023b) use explicit control signals to guide the temporal dynamics of the video. Gong et al. (Gong et al., 2023) introduce TaleCrafter to handle interactions among multiple characters, featuring layout and structural editing capabilities. He et al. (He et al., 2023) propose a retrieval-based deep guidance method that can integrate existing video clips into a coherent narrative video by customizing the appearance of characters. These studies mainly focus on the appearance of visual content or objects. Several methods (Zhao et al., 2023; Zhang et al., 2023; Wei et al., 2024; He et al., 2024) learn and controll motion through customized diffusion models. These attempts have made pioneering progress in controlling video in specific aspects. However, there is still a lack of effective solutions for precisely controlling the lighting effects in videos.
3 Methods
Our goal is to achieve unified control over video lighting, a challenging task with considerable implications. The primary difficulties are threefold: (1) Dataset Scarcity: there is a significant lack of lighting datasets, particularly for videos, with few annotated examples that explicitly showcase lighting changes and where lighting information is well-defined. (2) Complexity of Lighting Attributes: lighting encompasses various factors, including the type of light source, direction of illumination, and the material properties of objects. In the context of text-to-video generation, the material and shape of the generated objects are often unknown. Thus, effectively representing lighting information to accurately convey its visual effects within the camera’s field of view poses a substantial challenge. (3) Attribute Decoupling: like many control tasks, lighting control faces the challenge of decoupling specific attributes. A key technical hurdle is how to isolate lighting information from the appearance of the training data, ensuring that the model does not overfit to the training data’s appearance or incorporate irrelevant details.
Thus, this section is organized as follows. Section 3.1 presents the preliminary content of the video generation models; Section 3.2 introduces the method of collecting the lighting dataset that we propose, namely LumiHuman. Section 3.3 introduces the method of representing camera information and injecting it into the text-to-video model, namely LumiSculpt.
3.1 Preliminary
Text-to-video diffusion models
Recently, significant advancements have been made in Text-to-Video (T2V) diffusion models (Chen et al., 2023; Wang et al., 2023a; Lab & etc., 2024; Hong et al., 2022). Most of these models follow the classic algorithmic framework from the field of image generation. Specifically, the process typically involves gradually introducing noise $\epsilon$ into $N$ sequences of images $z_{1},\ldots,z_{N}$ until they approximate a Gaussian distribution. Given noise-corrupted inputs $z_{1},\ldots,z_{N}$, a neural network is trained to predict the added noise. During training, the network strives to minimize the mean squared error (MSE) between its noise predictions and the actual noise.
Figure 2: The collection process of LumiHuman includes: (a) designing a 3D point light source matrix, (b) rendering single-frame images and generating portrait lighting videos with various path lighting and lighting reference videos, (c) annotating with a BLIP model, and (d) producing enhanced background captions using a large language model.
3.2 LumiHuman
Lighting Videos
Extracting lighting information from existing videos presents significant challenges, even for professionals. The complexity of lighting encompasses factors such as light source type, angle, and distance. Consequently, annotating existing video data to create lighting datasets is challenging. Generating data with known lighting parameters through generative methods is a more viable approach.
We introduce a portrait lighting dataset, referred to as LumiHuman. This dataset is created using a unreal engine for lighting simulation, allowing for the production of data with known lighting information. The LumiHuman collection comprises five key stages: (1) Lighting Design: As illustrated in Figure 2(a), we developed a lighting position matrix, i.e., a three-dimensional grid measuring $160cm\times 160cm\times 160cm$. The points are uniformly spaced at $5cm$ intervals to serve as lighting positions. Point light sources move within these grid points to capture data of the subject being illuminated from various angles. (2) Lighting Trajectories Design: Within the three-dimensional grid, we defined horizontal, vertical, and diagonal trajectories, each composed of grid points to simulate diverse lighting change effects. (3) Character Construction: To facilitate the production of portrait lighting data, we utilized the MetaHuman dataset (MetaHuman, 2023), which features 3D models of 65 different individuals. This diversity enhances the visual effects of light projection across different characters. (4) Flexibility and Storage Optimization: To achieve a wide variety of lighting path variations while addressing storage concerns, i.e., given the presence of duplicate frames at the same lighting positions, we offer a flexible and lightweight image-video dataset. This dataset includes images rendered from various lighting positions in a three-dimensional grid, where each character corresponds to $33\times 33\times 3$ 33 different sampled images. In practice, videos can be generated from images in the dataset using predefined paths, or additional paths can be designed to simulate different lighting effects, as shown in Figure 2 (b). (5) Text Annotation and Augmentation: For automatic text annotation of the videos, we employed BLIP (Li et al., 2022). Additionally, we utilized GPT-4 (Achiam et al., 2023) for caption augmentation, generating diverse contextual backgrounds for the dataset, as shown in Figures 2(c) and (d).
Lighting Representation
To describe the effect of the lighting projection, a straightforward approach is to incorporate the lighting parameters as additional information in the model. However, this method requires a substantial amount of annotated data to establish a mapping between the lighting vectors and the two-dimensional plane. To better align with the model’s inferential feature space, we propose projecting lighting information into an empty space, as illustrated in Figure 3(b). For different lighting positions, this is represented as an image where brighter areas indicate stronger illumination, and darker areas signify weaker lighting. This representation allows for a more effective alignment of lighting information with the video generation model.
Figure 3: The pipeline of LumiSculpt consists of the generation backbone, i.e. the controlled branch, which includes (a) a diffusion transformer (DiT), a pre-trained video denoising network, and (b) a light encoder, a trainable external transformer network. The light encoder takes light reference latents as input and processes them through various blocks to produce a light condition sequence. This sequence is integrated into the generation backbone using several (d) merge modules within each block. During training, aside from the controlled branch, (c) a frozen branch is introduced to provide regularization for diverse appearances. The frozen branch is a DiT with frozen parameters, sharing weights with (a). Both branches predict noise, resulting in $\epsilon_{pred}$ and $\epsilon_{reg}$, which are used to compute the disentanglement loss $\mathcal{L}_{dis}$.
3.3 LumiScuplt
3.3.1 Integrating Lighting into Video Generators
Since lighting is represented in pixel space, it can be parameterized as input to the standard visual model. We extract lighting features using a Variational Autoencoder (VAE) shared with a generative model, which are then fed into a dedicated lighting encoder. As illustrated in Figure 3(b), this lighting encoder employs a transformer architecture comprised of self-attention layers, enabling it to compute attention scores globally across the video. This design allows the encoder to effectively capture the spatial and temporal relationships of lighting throughout the video clip. The lighting encoder accepts lighting features as input. The transformer blocks, matching the number of layers in the backbone model, output a sequence of latents of the same size. For each layer in the backbone model, the lighting encoder provides corresponding features of identical dimensions, facilitating feature fusion. Our objective is to seamlessly integrate these latents into the DiT architecture of the T2V model, as shown in Figure 3(d). The latent features of the video, denoted as $z_{t}$, and the lighting features, $c_{t}$, are combined through element-wise addition. This integrated feature is then passed through a linear layer, producing the output for the subsequent layer with a hyper-parameter termed guidance scale set to 0.5.
3.3.2 Lighting Learning
We utilize a data-driven approach to learn complex lighting information, capturing the projection effects of lighting at different positions on the human face. The implementation of this method is based on the LumiHuman. There is an issue with simply learning lighting from data rendered by unreal engines, which is the leakage of appearance information. This occurs due to the consistent backgrounds and layouts in the dataset. While such consistency provides more stable training data for lighting learning, it simultaneously increases the model’s susceptibility to overfitting to similar appearances. This raises the key issue of lighting control: how to decouple lighting from other elements? We propose a lighting-appearance disentanglement method, which includes a dual-branch framework and a novel disentanglement loss.
Dual-Branch Framework
To obtain diverse appearance data, an naive solution is to use additional video data as regular samples to provide the model with diverse appearances. However, it is challenging to obtain a large amount of diverse data with known lighting conditions, so we opt for the manufacture of regular samples based on generative models. As shown in Figure 3(c), we propose a dual-branch structure, including a training branch and a frozen branch, by introducing a frozen foundational denoising model to provide an appearance reference. During training, the dual branches accept the same textual conditions and noisy latents, obtaining predicted noises $\epsilon_{t}$ and $\epsilon_{t}^{reg}$, respectively. The diverse appearances of the pre-trained model are reflected in $\epsilon_{t}^{reg}$. In this way, we achieve low-cost regular sample manufacturing.
Loss Functions
We train the model to learn the overall distribution of the dataset using a simple denoising loss $\mathcal{L}_{denoise}$, which includes lighting information and appearance information from the training videos, the latter of which we do not require. To remove appearance, a quantitative method (e.g., a loss function) capable of measuring appearance consistency without affecting lighting learning is needed. On the one hand, this loss should be able to quantify the appearance consistency of the latents from two videos, and on the other hand, the loss must be independent of planar location, otherwise it will affect the learning of lighting distribution. Inspired by style transfer (Jing et al., 2020), which also needs to measure style consistency without being disturbed by structured information, we propose a disentanglement loss $\mathcal{L}_{dis}$ to measure appearance consistency. We use AdaIN (Huang & Belongie, 2017) to measure the feature distribution patterns of two videos’ latents, reflecting their appearance feature distribution. Aligning the feature distribution of the training branch-generated video with the frozen branch retains the diverse appearance of the original model. With this approach, the generative model globally aligns with the training video under the drive of $\mathcal{L}_{denoise}$, while squeezing out redundant appearance information under the drive of $\mathcal{L}_{dis}$, retaining the generative capability of the original model. The training objective function can be represented as:
$\mathcal{L}_{dis}$ $=||(\sigma(z^{pred}_{0})-\sigma(z^{reg}_{0})||_{2}+||(\mu(z^{pred}_{0})-\mu(z^{reg}_{0})||_{2},$ (1)
$\mathcal{L}_{denoise}$ $=\mathbb{E}_{z_{1:N},\epsilon,c_{t},t}\left[||\hat{\epsilon}(z^{pred}_{1:N},c_{t},t)-\epsilon||^{2}\right],$
$\mathcal{L}_{total}$ $=\mathcal{L}_{denoise}+\beta\mathcal{L}_{dis},$
where $\sigma(\cdot)$ and $\mu(\cdot)$ denote the computation of standard variance and mean, respectively. $z_{0}^{reg}=z_{t}^{reg}-\epsilon_{reg}$ represents the predicted denoised output of the frozen branch at time step $t$, while $z_{0}^{pred}=z_{t}^{pred}-\epsilon_{pred}$ signifies the predicted denoised output of the controlled branch at time step $t$. $N$ denotes the total number of steps, and $c_{t}$ represents the textual condition. $\beta$ is set to 3.0.
4 Experiments
4.1 Experimental Setup
Methods for comparison  We compare our approach with state-of-the-art text-to-video generation methods Open-Sora (Lab & etc., 2024) and relighting method IC-light (Zhang et al., 2024).
Metrics  We employ a variety of quantitative and qualitative metrics to assess the lighting accuracy, inter-frame coherence, and visual-text similarity of generated videos.
Evaluation dataset  We use 500 different light paths and captions not present in the training dataset as conditions to guide the comparative methods in generating evaluation videos.
Implementation details. In all video generation experiments, we use Open-Sora v1.2.0 (Lab & etc., 2024) with the default network architecture. We set a learning rate of $1\times 10^{-4}$. The input video resolution is $640\times 480\times 29$. The training process for each motion requires approximately $800\sim 1500$ iterations using eight NVIDIA A100. The number of inference steps is set to $T=50$ and the guidance scale is set to $w=7.5$.
Table 1: Quantitative experimental results and ablation study results.The best results are marked as bold and the seconds one are marked by underline.
Method Consistency Lighting Accuracy Quality
[…]
Open-Sora 0.9845 1.3503 0.4542 0.8229 0.3182
IC-Light 0.9703 2.5329 0.5264 0.8632 0.3145
Ours(full model) 0.9951 1.1312 0.3500 0.8779 0.3597
Ours(w/o caption aug) 0.9948 1.1211 0.2992 0.9269 0.3416
Ours(w/o $\mathcal{L}_{dis}$) 0.9957 1.1033 0.1945 0.8363 0.2909
Figure 4: Comparison results with state-of-the-art methods IC-Light (Zhang et al., 2024) and Open-Sora (Lab & etc., 2024). The classic horizontal and vertical directions for light movement and two brightness levels are tested to achieve a comprehensive qualitative evaluation.
4.2 Quantitative Evaluations
As shown in Table 1, we use five quantitative metrics for evaluation: (1) Frame-wise CLIP image similarity: we utilize the similarity of frame-wise CLIP (Radford et al., 2021) image embeddings to evaluate the semantic-level video coherence. A higher value indicates greater inter-frame similarity, suggesting better semantic stability in the generated video. (2) Frame-wise Learned Perceptual Image Patch Similarity (LPIPS) consistency: we measure feature-level coherence using frame-wise LPIPS consistency. A lower value signifies smaller feature discrepancies, indicating higher inter-frame consistency. (3) Light directions Root Mean Squared Error (RMSE): we calculate the lighting direction for each frame and then assess the consistency of the generated video’s lighting direction with the reference. RMSE represents the discrepancy in lighting direction; the smaller the value, the more consistent the lighting with the target. (4) Brightness consistency: we segment each video frame into patches, compute the average brightness of different patches, and construct a brightness distribution relationship. This distribution is only related to the comparison between different patches and is independent of the absolute brightness values. We calculate the brightness distribution consistency between the generated video and the reference video. (5) CLIP text-image similarity: we measure the model generation quality using the similarity between the clip image embedding of the generated video frames and the text embedding of the caption. The higher the similarity, the better the generation quality. It can be observed that, compared to Open-Sora and IC-Light, LumiSculpt is capable of maintaining good inter-frame consistency and text-image consistency while achieving accurate lighting control.
4.3 Qualitative Evaluations
As shown in Figure 4, due to the absence of video illumination control methods, we compare our approach with the image illumination control method IC-Light based on diffusion models, and the video generation method Open-Sora. We consider two light intensity levels, strong and soft, as well as horizontal and vertical lighting movement directions. Since IC-Light is designed for relighting existing images, we use portraits generated by our method as foreground guidance. IC-Light is capable of producing single-frame images with accurate lighting directions, but due to a lack of inter-frame awareness, the coherence of the output video is poor, with noticeable flickering in the background. Open-Sora can generate coherent and aesthetically pleasing videos, but struggles to control lighting direction via textual conditions, resulting in relatively unchanged lighting throughout the video. Our method not only ensures video coherence and visual quality but also achieves precise control over lighting trajectory and intensity. Video results are provided in the supplementary material.
Figure 5: Ablation study results. We conducted ablations on key components related to model training, including caption augmentation and $\mathcal{L}_{dis}$, both of which contribute to the diversity of generated content. We presented results with varying values of the important parameter guidance scale, which affects the accuracy of lighting and the diversity of appearance, related to model inference.
Table 2: Experimental results of hyper-parameter guidance scale. The best results are marked as bold and the seconds one are marked by underline.
[…]
scale=0.1 0.9943 0.4239 0.3163
scale=0.3 0.9964 0.3825 0.3814
scale=0.5 0.9951 0.3500 0.3597
scale=0.7 0.9939 0.2484 0.3499
scale=0.9 0.9902 0.2922 0.2941
4.4 Ablation Study
As shown in the $1^{st}\sim 3^{rd}$ and last rows of Figure 5, we present the results of ablating different modules of LumiSculpt. Removing caption augmentation from LumiHuman leads to a lack of diverse textual guidance, causing the model during training to rely solely on text conditions that exactly match the dataset, thus improving appearance fitting. As shown in the second row, the generated results exhibit consistent pose and layout. Without the dual-branch structure and decoupling loss, as shown in the third row, the generated appearances tend to overfit the training data, making it challenging to produce diverse backgrounds. As illustrated in the first row, the complete LumiSculpt successfully balances diverse appearances with accurate lighting. As shown in the $4^{th}\sim 7^{th}$ rows of Figure 5 and Table 2, we present both qualitative and quantitative analysis results of varying the hyper-parameter guidance scale. During inference, the standard LumiSculpt sets the guidance scale to 0.5. Increasing the guidance scale intensifies the strength of lighting guidance, enhancing the accuracy of lighting direction and brightness, but an excessively large guidance scale can undermine the model’s ability to generate diverse outputs. In contrast, decreasing the guidance scale reduces the strength of lighting guidance, leading to a decrease in lighting accuracy.
5 Conclusion
In this paper, we address the challenge of lighting control in text-to-video generation. To address data scarcity, lighting representation difficulties, and lighting injection complexities, we introduce a flexible portrait lighting dataset, LumiHuman, along with a plug-and-play lighting guidance method, LumiSculpt. Our proposed dual-branch structure and associated loss function for decoupling are not only effective for lighting control but also have the potential to be generalized across a variety of generative tasks. As video generation techniques continue to evolve, enabling precise lighting control becomes an increasingly important research direction, driven by significant aesthetic demands from both professionals and the general public. We believe that LumiHuman and LumiSculpt will serve as valuable resources and methodologies for future explorations in this domain.